A car manufacturing company was able to release 1,860 units at the end of January. Based on records, they were able to increase their sales by 46 units per month thereafter. In what month did their car sales reach 2, 320 units?


1
Expert's answer
2021-12-01T17:58:53-0500

The persentage increase is:

46/1,860 = 0.0247 or 2.47%.

So, their car sales will reach 2,320 units in:

"1,860\u00d71.0247^x = 2,320,"

"1.0247^x = 1.2473,"

x = 9.06 month.

So, the car sales will reach 2,320 units in November.